Circular No. 7263

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SAX J1719.6–4254, MXB 1658–298, and SAX J1819.3–2525 J. Heise, Space Research Organization Netherlands, Utrecht; L. Na-

talucci and M. Cocchi, Istituto di Astrofisica Spaziale, CNR, Rome; and M. Smith, G. Tassone, G. Celidonio, and L. Di Ciolo, BeppoSAX Science Operations Center, Rome, communicate: “During an observation with Wide Field Camera 2 (WFC2) on BeppoSAX between Sept. 22.04 and 23.45 UT, a new faint x-ray transient was detected with a maximum flux of 23 mCrab (2–9 keV). The source, designated SAX J1719.6–4254, is located at α = 17^{h}19^{m}38^{s}, δ = −42^{o}53′.8 (equinox 2000.0), with a radius of the 99-percent-confidence error circle of 2′.1 The nearest SIMBAD source is a K2 star (V = 9.41) located 3′.3 away. In the same field-of-view of the WFC, MXB 1658–298 is visible at a flux level of ∼ 30 mCrab; the persistent emission of this source was not detected previously with the WFCs, but recently an x-ray burst was seen on Sept. 2.10. SAX J1819.3–2525 (GM Sgr; cf. IAUC 7253) was not detected above a (5σ) upper limit of 8 mCrab. A follow-up observation on SAX J1719.6–4254 with the BeppoSAX Narrow Field Instruments is being scheduled.”

RW URSAE MINORIS A. Retter, Keele University, reports that CCD photometry of the clas-

sical nova RW UMi (cf. IAUC 4672) during 14 nights in 1997, with the 1-m Wise telescope through a ‘clear’ filter, yields periodic modulations in the light curve with a peak-to-trough amplitude of 0.03 mag and period 0.0591 ± 0.0003 day.

DD CIRCINI Visual magnitude estimates communicated via H. C. Vital, Rio de

Janeiro, Brazil: Sept. 6.006 UT, 10.4 (E. Salazar, Cusco, Peru); 8.000, 10.6 (Salazar); 9.097, 10.8 (J. Garcia, Rama Caida, Argentina); 10.993, 10.9 (A. Alves, Florianopolis, Brazil); 14.194, 11.3 (Garcia); 15.937, 11.4 (Alves); 16.958, 11.5 (Alves); 18.951, 11.6 (Vital); 20.014, 11.8 (Garcia).

SUPERNOVA 1999dn IN NGC 7714 Unfiltered CCD magnitudes: Sept. 4.666 UT, 15.7 (S. Yoshida and K.

Kadota, Ageo, Saitama, Japan); 11.010, 16.3 (E. Prosperi, Larciano, Italy).

1999 September 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green